# Mechanisms for producing Primordial Black Holes from Inflationary Models Beyond Fine-Tuning

### v1

@@ 7 Conclusions
This study has provided an extensive evaluation of fine-tuning requirements across various inflationary models for the production of PBHs. We have demonstrated that while single field models with inflection points require significant fine-tuning to align with CMB observations and produce PBHs, alternative models with step-like features in the scalar potential present a more favourable scenario. These models significantly reduce the need for fine-tuning and are capable of explaining the potential low-frequency GW signals detectable by PTA collaborations.
Further exploration into multi-field models reveals that the addition of extra fields can considerably alleviate the fine-tuning demands. Specifically, models incorporating non-canonical kinetic terms and hybrid inflationary models with a waterfall trajectory have shown promising results, reducing the fine-tuning requirements by orders of magnitude compared to traditional single field models.
Moreover, the implementation of a spectator field offers a novel approach to circumvent the fine-tuning issue entirely. These models allow for the natural stochastic variations in field values across different Hubble patches, which can lead to significant curvature fluctuations necessary for PBH production without the stringent constraints on the inflationary potential’s parameters.
In this work we study inflationary models in order to understand the PBH formation. These insights help expand the theoretical framework connecting inflation, dark matter, and GWs. While contributing to cosmology, this study also suggests more avenues for refining these models to enhance our understanding of their observable signatures in GWs. We acknowledge the preliminary nature of these findings and the need for further exploration to fully realize their implications.
Appendix A Appendices
A.1 Evaluation of the scalar power spectrum
In this appendix we summarise the evaluation of power spectrum which is needed in our analysis. In this framework, the field equations for ${\varphi^{i}}$, expressed in terms of efold time (assuming $\rm{M_{P}=1}$), are provided by the following equation:
@@ A.1 Evaluation of the scalar power spectrum
where the dots indicate derivatives with respect to efold time, and the indices $\mathrm{i=a,b,c}$ denote derivatives relative to the fields. The Christoffel symbols $\Gamma^{c}_{ab}$ are defined as:
@@
where $\mathrm{G_{ij}}$ represents the field metric. The velocity field $\dot{\sigma}$ is calculated as:
@@
and the slow-roll parameter $\epsilon_{1}$ is determined from:
$\epsilon_{1}=\frac{1}{2}\dot{\sigma^{2}}.$ (28)
@@
$\begin{split}&\delta\ddot{\varphi}^{c}+(3-\epsilon_{1})\delta\dot{\varphi}^{c}+2\Gamma^{c}_{ab}\dot{\varphi}^{a}\delta\dot{\varphi}^{b}+\left(\Gamma^{c}_{ab,d}\dot{\varphi}^{a}\dot{\varphi}^{b}+\frac{V^{c}_{d}}{H^{2}}-G^{ca}G_{ab,d}\frac{V^{b}}{H^{2}}\right)\delta\varphi^{d}\\ &+\frac{k^{2}}{a^{2}H^{2}}\delta\varphi^{c}=4\dot{\Psi}\dot{\varphi}-2\Psi\frac{V^{c}}{H^{2}}\end{split}$ (29)
@@
$\ddot{\Psi}+(7-\epsilon_{1})\dot{\Psi}+\left(2\frac{V}{H^{2}}+\frac{k^{2}}{a^{2}H^{2}}\right)\Psi=-\frac{V_{c}}{H^{2}}\delta\varphi^{c}.$ (30)
We assume the initial condition of a Bunch-Davies vacuum. The power spectrum is described by:
@@
A.2 Evaluation of the fractional abundances of PBHs
In this appendix we briefly discuss how to evalaute the the fractional abundance of PBHs. The present abundance of PBH is given by the integral:
@@ A.2 Evaluation of the fractional abundances of PBHs
Here $\gamma$ is a factor which depends on gravitation collapse and we choose $\gamma=0.2$ [85]. $T_{f}$ denotes the temperature of PBH formation and $M_{\text{PBH}}$ is the mass of PBHs. $g(T_{f})$ are the effective degrees of freedom during this formation and counting only the SM particles we set $g(T_{f})=106.75$.
The mass fraction of PBH at formation, denoted by $\beta(M_{\text{PBH}})$:
@@
where $\delta$ denotes
@@
The mass as a function of $k$ [24]:
@@
A.3 Energy Density of GWs
In this appendix we briefly discuss how to evaluate the energy density of GWs from the scalar power spectrum. The energy density of the GWs can be evaluated from the following integral  [86, 87, 88, 89, 90]
${\Omega_{GW}(k)}=\frac{\Omega_{r}}{36}\int^{\frac{1}{\sqrt{3}}}_{0}\mathrm{d}d\int^{\infty}_{\frac{1}{\sqrt{3}}}\mathrm{d}s\left[\frac{(s^{2}-1/3)(d^{2}-1/3)}{s^{2}+d^{2}}\right]^{2}\,P_{R}(kx)\,P_{R}(ky)\,(I_{c}^{2}+I_{s}^{2}).$ (38)
@@ A.3 Energy Density of GWs
we can evaluate the energy density of power spectrum as a function of the frequency $f$.

### v2

@@ 7 Conclusions
This study has provided an extensive evaluation of fine-tuning requirements across various inflationary models for the production of PBHs. We have demonstrated that while single-field models with inflection points require significant fine-tuning to align with CMB observations and produce PBHs, alternative models with step-like features in the scalar potential present a more favourable scenario. These models significantly reduce the need for fine-tuning and are capable of explaining the potential low-frequency GW signals detectable by PTA collaborations.
Further exploration into multi-field models reveals that the addition of extra fields can considerably alleviate the fine-tuning demands. Specifically, models incorporating non-canonical kinetic terms and hybrid inflationary models with a waterfall trajectory have shown promising results, reducing the fine-tuning requirements by orders of magnitude compared to single-field models.
Moreover, the implementation of a spectator field offers a novel approach to circumvent the fine-tuning issue entirely. These models allow for the natural stochastic variations in field values across different Hubble patches, which can lead to the significant curvature fluctuations necessary for PBH production without stringent constraints on the inflationary potential’s parameters.
To sum up, we examined several inflationary models for the production of PBHs and classified them based on the degree of fine-tuning required for explaining their copious production. We provided a comparative analysis of single- and multi-field inflationary mechanisms, highlighting the advantages of multi-field approaches in reducing the need for fine-tuning. These insights help expand the theoretical framework connecting inflation, PBHs, and GWs. The novelty of this study lies in its contribution to guiding future research toward models that offer a more natural explanation for the formation of PBHs, thus paving the way for more grounded investigations in this area.
[…]
The work of I.S. from the Belgian Francqui Foundation through a Francqui Start-up Grant, as well as the Belgian Fund for Scientific Research through an MIS grant and an IISN grant.
[…]
No new data were created or analyzed in this study. Data sharing is not applicable to this article.
[…]
The authors declare no conflicts of interest.
[…]
Appendix A
In this Appendix A, we summarize the evaluation of the power spectrum that is needed in our analysis. Additionally, we outline several noteworthy characteristics of the models explored in this study.
In this framework, the field equations for ${\varphi^{i}}$, expressed in terms of efold time (assuming ${M_{P}=1}$), are provided by the following equation:
@@
where the dots indicate derivatives with respect to efold time, and the indices ${i=a,b,c}$ denote derivatives relative to the fields. The Christoffel symbols $\Gamma^{c}_{ab}$ are defined as:
@@
where ${G_{ij}}$ represents the field metric. The velocity field $\dot{\sigma}$ is calculated as:
@@
and the slow-roll parameter $\varepsilon_{1}$ is determined from:
$\varepsilon_{1}=\frac{1}{2}\dot{\sigma^{2}}.$ (28)
@@
$\begin{split}&\delta\ddot{\varphi}^{c}+(3-\varepsilon_{1})\delta\dot{\varphi}^{c}+2\Gamma^{c}_{ab}\dot{\varphi}^{a}\delta\dot{\varphi}^{b}+\left(\Gamma^{c}_{ab,d}\dot{\varphi}^{a}\dot{\varphi}^{b}+\frac{V^{c}_{d}}{H^{2}}-G^{ca}G_{ab,d}\frac{V^{b}}{H^{2}}\right)\delta\varphi^{d}\\ &+\frac{k^{2}}{a^{2}H^{2}}\delta\varphi^{c}=4\dot{\Psi}\dot{\varphi}-2\Psi\frac{V^{c}}{H^{2}}\end{split}$ (29)
@@
$\ddot{\Psi}+(7-\varepsilon_{1})\dot{\Psi}+\left(2\frac{V}{H^{2}}+\frac{k^{2}}{a^{2}H^{2}}\right)\Psi=-\frac{V_{c}}{H^{2}}\delta\varphi^{c}.$ (30)
We assume the initial condition of a Bunch–Davies vacuum. The power spectrum is described by:
@@
In the case of single-field inflation $\varphi^{1}=\phi$, the equation of the curvature perturbations takes the following form Kefala et al. (2021):
${R^{\prime\prime}}_{k}+f(N)R^{\prime}_{k}+\frac{k^{2}}{H^{2}}e^{-2N}R_{k}=0$ (33)
where
$f(N)=3+2\frac{\phi^{\prime\prime}}{\phi^{\prime}}-\frac{\phi^{\prime 2}}{2}\,.$ (34)
This function acts as a friction term.
The potential in Equation (8) can interplay between the mechanism of the inflection point and that of the step-like potential. In Figure 7, we illustrate the behavior of the slow-roll parameter $\varepsilon_{1}$, Equation (28), and the $f$ from Equation (34) for two different scenarios: an inflection point (blue curve) and a step-like potential (orange line), using the potential specified in Equation (8). The parameters for each scenario are provided in the figure. This plot highlights the distinct dynamics between these two mechanisms; in the inflection point scenario, the velocity of the field significantly decreases, whereas in the step-like potential, the field’s velocity exhibits both increases and decreases. Moreover, the function $f$ behaves differently in the cases of the inflection point and step-like potential. Hence, through this figure we can notice that the previously mentioned mechanisms behave differently.
Figure 7: The different behavior of the slow-roll parameter $\varepsilon_{1}$ (left panel) and friction term $f$ (right panel) for the mechanism of an inflection point (blue line) and the one with the step-like potential (orange line). For blue line: $i=1$, $A_{1}=0.000605$, $c=100$, $B=-0.03$. For orange line: $i=3$, $A_{1}=0.1$, $c=100$ $B=0.03$, $\varphi_{1}=0.4$, and $\varphi 2=0.8$. For initial condition, $\phi_{ic}=-0.3$.
Appendix B
In Appendix B, we briefly discuss the methodology to evaluate the fractional abundance of PBHs. First, we outline the theoretical framework and computational approaches used to estimate the PBH abundance. Subsequently, we present the fractional abundances calculated for the various inflationary models discussed within this work. Finally, we discuss these results.
The fractional abundance of PBH is given by the integral:
@@
Here, $\gamma$ is a factor that depends on gravitation collapse, and we choose $\gamma=0.2$ Carr (1975). $T_{f}$ denotes the temperature of PBH formation, and $M_{\text{PBH}}$ is the mass of PBHs. $g(T_{f})$ are the effective degrees of freedom during this formation, and counting only the SM particles we set $g(T_{f})=106.75$.
The mass fraction of PBHs at formation, denoted by $\beta(M_{\text{PBH}})$, can be estimated using either the Press–Schechter (PS) approach Press and Schechter (1974) or Peak Theory Bardeen et al. (1986). In this section, we focus on the PS approach. Hence, the mass fraction of $\beta(M_{\text{PBH}})$ is given as:
@@
where the overdensity $\delta$ is above a certain threshold of collapse $\delta_{c}$. The variance of curvature perturbation is given :
@@
The mass as a function of the comoving wavenumber $k$ Ballesteros and Taoso (2018):
@@
Figure 8: The fractional abundances of PBHs for the four different mechanisms: inflection point (blue), step-like feauture potential (red), two field with a non-canonical kinetic term (orange), and hybrid model (green).
In Figure 8, we illustrate the fractional abundances of PBHs for various mechanisms: an inflection point, step-like features in the potential, a two-field model with non-canonical kinetic terms in the Lagrangian, and a two-field model with a waterfall trajectory, each represented by distinct colors as detailed in the caption. We use the parameter sets presented in Figures 14, which lead to an enhancement of the scalar power spectrum of at least $\rm{10^{-2}}$. We can notice that models with step-like potential can lead to narrower mass ranges. The models shown in Figure 8 can account for a wide range of PBH masses. However, it is important to emphasize that different parameter choices can yield, varying fractional abundances; therefore, this plot should not be solely relied upon for comparing these mechanisms. Furthermore, the fractional abundances can exceed constraints, such as those given in Refs. Capela et al. (2013); Niikura et al. (2019). As mentioned, the proper adjustment of parameters for each model is required, but this is beyond the scope of this paper.
For the evaluation of fractional abundance depicted in Figure 8, we adopt critical threshold values ranging from $\delta_{c}=0.4$ to $0.6$ as referenced in Ref. Musco et al. (2021); Escrivà et al. (2021, 2022). It is crucial to compute the threshold value from the respective power spectrum for each scenario, as outlined in Musco et al. (2021); Escrivà et al. (2021, 2022); Stamou (2023). Nonetheless, by appropriately adjusting the model parameters, one can achieve the desired fractional abundances. The choice of window function, specifically a Gaussian window function as applied in this analysis (see Equation (38)), significantly influences the threshold calculation. Although the choice of window function introduces corrections in the threshold computation, these impacts are minimized when the same window function is consistently used across evaluations of PBH abundance Kalaja et al. (2019); Young (2019); Tokeshi et al. (2020); Musco et al. (2021).
Appendix C
In this Appendix, we briefly discuss how to evaluate the energy density of GWs from the scalar power spectrum. The stochastic background of GWs can be generated through the coupling of the scalar and tensor modes Espinosa et al. (2018); Kohri and Terada (2018); Mollerach et al. (2004); Maggiore (2000); Baumann et al. (2007).
The energy density of the GWs can be evaluated from the following integral:
${\Omega_{GW}(k)}=\frac{\Omega_{r}}{36}\int^{\frac{1}{\sqrt{3}}}_{0}\mathrm{d}d\int^{\infty}_{\frac{1}{\sqrt{3}}}\mathrm{d}s\left[\frac{(s^{2}-1/3)(d^{2}-1/3)}{s^{2}+d^{2}}\right]^{2}\,P_{R}(kx)\,P_{R}(ky)\,(I_{c}^{2}+I_{s}^{2})\,,$ (40)
@@
we can evaluate the energy density of the power spectrum as a function of the frequency $\rm{f}$.
{adjustwidth}
[…]